OpenStudy (anonymous):

Explain how you would know log3100 is between 4 and 5 without the use of a calculator.

OpenStudy (anonymous):

log3 (100) = x so 3^x = 100 What value of x would give you 3^x = 100? 3^2 = 9, 3^3 = 27, 3^4 = 81, 3^5 = 243 100 is between 81 and 243 so x must be between 4 and 5.

OpenStudy (anonymous):

\[\log_{3} 100\]

OpenStudy (anonymous):

change it into index form\[3^x=100\]

OpenStudy (anonymous):

thanks guys! yea i got to 81 and 243 and then got confused and thought i did it wrong :)

OpenStudy (anonymous):

how do i mark that is question has a good answer?

OpenStudy (anonymous):

Click on Good Answer for one of the given answers...

OpenStudy (anonymous):

do u mean the quwstion or the answer as a good answer

OpenStudy (anonymous):

ugh my computer is freezing and typing weird sorry